# Citadel: Context Similarity Based Deep Learning Framework Bug Finding

### v4

@@ 8. Conclusion
This paper presents Citadel, a bug-finding tool for DL frameworks that can find new bugs that are similar to known bugs, regardless of bug types. For a problematic DL framework API function and its associated bugs, it matches analogous API functions from the perspectives of context similarity and signature similarity and then synthesizes test cases for these analogous API functions. Then, Citadel leverages the behavior of the confirmed bug on the problematic API as the test oracle to evaluate the generated test cases and identify new bugs on analogous API functions. Our evaluation on two frameworks shows that Citadel can effectively and efficiently detect status, value, and performance bugs.

### v5

@@ 8. Conclusion
This paper presents Citadel, a bug-finding tool for DL frameworks that can find new bugs that are similar to known bugs, regardless of bug types. For a problematic DL framework API function and its associated bugs, it matches analogous API functions from the perspectives of context similarity and signature similarity, and then synthesizes test cases for these analogous API functions. Then, Citadel leverages the behavior of the confirmed bug on the problematic API as the test oracle to evaluate the generated test cases and efficiently identify new API bugs on analogous API functions. Our evaluation on two frameworks shows that Citadel can effectively and efficiently detect status, value, and performance bugs.
Acknowledgements.
The authors would like to thank the anonymous reviewers for their insightful comments and valuable suggestions. Authors in China are supported partially by the National Key Research and Development Program of China (2023YFE0209800), the National Natural Science Foundation of China (U24B20185, T2442014, 62161160337, 62132011, U21B2018), and the Shaanxi Province Key Industry Innovation Program (2023-ZDLGY-38, 2021ZDLGY01-02). Thanks to the New Cornerstone Science Foundation and the Xplorer Prize.